Copy Machine Use A store manager hypothesizes that the average number of pages a person copies on the store’s copy machine is less than 40. A sample of 50 customers’ orders is selected. At a _ 0.01, is there enough evidence to support the claim? Assume population standard deviation 30.9.




1
Expert's answer
2022-01-06T11:46:19-0500

"n=50"

"s=30.9"

"\mu_0=40"

X = Number of pages a person copies on the store's copy machine

"\bar{X}=\frac{\Sigma X}{n}=\frac{1463}{50}=29.26"


Null & Alternative Hypothesis:

"H_0:\mu=\mu_0"

"H_a:\mu<\mu_0"

Test Statistic:

Sample size is greater than 30, so Z-test is appropriate to test null hypothesis.

"Z=\frac{\bar{X}-\mu}{\frac{\bar{s}}{\sqrt{n}}}"

"Z=\frac{29.26-40}{\frac{\bar{30.9}}{\sqrt{50}}}=-2.46"

Critical Value:

The level of significance is 0.01.

"Z_\alpha=Z_0.01=-2.326"

Statistical Decision:

Since, Z-test statistic is less than critical value (-2.46 < -2.326). So, the null hypothesis is rejected.

Conclusion:

It can conclude that average number of pages a person copies on the store's copy machine is less than 40.